a man has $16000 which he places at interest,part at 3 1/2 percent the rest at 4 percent.his total annual income from this investment is $597.50.how much is invested at each rate?
Answer by mananth(16946)

Part I 3.50% per annum ------------- Amount invested =x
Part II 4.00% per annum ------------ Amount invested = y
16000
Interest----- 597.50

Part I 3.50% per annum ---x
Part II 4.00% per annum ---y
Total investment
x + 1 y= 16000 -------------1
Interest on both investments
3.50% x + 4.00% y= 597.5
Multiply by 100
3.5 x + 4 y= 59750.00 --------2
Multiply (1) by -3.5
we get
-3.5 x -3.5 y= -56000.00
Add this to (2)
0 x 0.5 y= 3750
divide by 0.5
y = 7500
Part I 3.50% $ 8500
Part II 4.00% $ 7500

CHECK
8500 --------- 3.50% ------- 297.50
7500 ------------- 4.00% ------- 300.00
Total -------------------- 597.50

m.ananth@hotmail.ca